In a psychology class there are 13 men and 12 women. 2 students are randomly selected to present a topic. What is the probability that at least 1 of the 2 students selected is male? Express your answer as a fraction or a decimal number rounded to four decimal places.
Answer #1

Answer

total number of students = 12+13 = 25 students

Number of ways of selecting 2 male out of 13 males = combination(13,2) = 13!/[(13-2)!*2!] = 13!/[11!*2!] = 78 ways of selecting both students as men

Number of ways of selecting at 1 male and 1 female = combination(13,1)*combination(12,1)

this gives

combination(13,1)*combination(12,1) = (131/|(13-1 ) ! * 11) * ( 12!Д( 12-1 ) ! * 11) 156ways

And number of ways of selecting 2 students out of 25 = combination(25,2) = (25!/[(25-2)!*2!]) = 25!/[23!*2!] = 300 ways

So, the required probability of selecting at least 1 male = (number of ways of 2 male+number of ways of selecting 1 male and 1 female)/(total number of ways of selecting 2 students out of 25 students)

setting the values, we get

Probability = (78+156)/300 = 39/50 = 0.7800

So, the required probability is 39/50